# Model Collapse in the Self-Consuming Chain of Diffusion Finetuning: A Novel Perspective from Quantitative Trait Modeling

### v2

@@ Abstract
The success of generative models has reached a unique threshold where their outputs are indistinguishable from real data, leading to the inevitable contamination of future data collection pipelines with synthetic data. While their potential to generate infinite samples initially offers promise for reducing data collection costs and addressing challenges in data-scarce fields, the severe degradation in performance has been observed when iterative loops of training and generation occur—known as “model collapse.” This paper explores a practical scenario in which a pretrained text-to-image diffusion model is finetuned using synthetic images generated from a previous iteration, a process we refer to as the “Chain of Diffusion.” We first demonstrate the significant degradation in image quality caused by this iterative process and identify the key factor driving this decline through rigorous empirical investigations. Drawing an analogy between the Chain of Diffusion and biological evolution, we then introduce a novel theoretical analysis based on quantitative trait modeling. Our theoretical analysis aligns with empirical observations of the generated images in the Chain of Diffusion. Finally, we propose Reusable Diffusion Finetuning (ReDiFine), a simple yet effective strategy inspired by genetic mutations. ReDiFine mitigates model collapse without requiring any hyperparameter tuning, making it a plug-and-play solution for reusable image generation.
@@ 1 Introduction
Can state-of-the-art AI models learn from their own outputs and improve themselves? As generative AI models (e.g., GPT, Diffusion) now churn out uncountable synthetic texts and images, this question piqued the curiosity of many researchers in the past couple of years. While some show positive results of self-improving [1, 2], most report an undesirable “model collapse”—a phenomenon where a model’s performance degrades when it undergoes multiple cycles of training with the self-generated data [3, 4, 5, 6, 7, 8, 9, 10]. When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text with many repetitions [11], and its linguistic diversity declines rapidly [12, 13]; image models also show quality degradation [14, 15] and loss of diversity [16, 17].
@@
The goal of this paper is to investigate model collapse in the practical scenario of finetuning pretrained text-to-image diffusion models. An end user of text-to-image diffusion models often wants to finetune the latest model to generate images with a very specific style (e.g., creating new characters in the style of Pokémon). In fact, hundreds of new finetuned diffusion models are regularly uploaded on platforms like CivitAI [^1] https://civitai.com/, each designed to produce different styles of images. When users scrape the internet to collect images in the style they want, it becomes almost inevitable that synthetic images will be included in their datasets. This is because the number of real images is limited, while synthetic images can be generated in massive quantities and dominate online sources. As a result, users will feel compelled to include more synthetic images in their datasets to meet the data demands of these ever-growing, data-hungry models.
@@
To this end, we conduct a thorough investigation into how various hyperparameters commonly used during diffusion finetuning (e.g., learning rate, diffusion steps, prompts) impact model collapse. From our extensive empirical analysis, we make a crucial observation: the classifier-free guidance (CFG) scale is the most significant factor influencing the rate of model collapse. Moreover, we observe a fascinating phenomenon that the direction of degradation varies with different CFG values—low CFG results in low-frequency degradation in images, while high CFG leads to high-frequency degradation. The critical role of CFG in determining both the rate and direction of model collapse is a novel insight previously unrecognized in the literature.
@@
To gain a deeper theoretical understanding of this phenomenon, we draw upon the concept of quantitative trait modeling from statistical genetics. Unlike existing studies that attribute model collapse to limited sample sizes and conclude that it leads to zero variance [16, 3, 6], we propose that the underlying cause is a truncation-based selection process modulated by CFG. This theoretical model accurately describes our experimental observations. Moreover, we show that other theoretical work on model collapse can also be connected to statistical genetics models, such as random genetic drift and its variants [18, 19, 20, 21]. This fresh angle of drawing parallels with statistical genetics offers a new framework to understand the mechanism of model collapse.
Finally, inspired by our empirical results and theoretical insights, we introduce the Reusable Diffusion Finetuning (ReDiFine) method. Our experiments show that selecting an appropriate CFG can significantly slow down model collapse. However, finding the optimal CFG is computationally expensive, requiring numerous iterations of finetuning. To address this, we propose ReDiFine as a plug-and-play solution that achieves a comparable reduction in collapse rate without the need for any hyperparameter tuning. The default ReDiFine setting effectively mitigates model collapse across all four datasets we tested. Moreover, ReDiFine is effective at mitigating model collapse not only in a worst-case scenario of having a fully synthetic dataset, but also in a more practical case, where we have a mix of real and synthetic images in the training set. Our contributions can be summarized as follows:
  • We performed thorough and extensive empirical investigations into model collapse when we finetune a diffusion model with diffusion-generated images. We tested an exhaustive list of experimental parameters on four datasets (two digital art & two natural images) to identify what affects model collapse. Our analysis reveals that CFG scale is the most influential factor that not only controls the rate of model collapse but also dictates the type of image degradation (Section 3).
@@
  • We propose a simple yet effective strategy to mitigate model collapse, named ReDiFine, which combines condition drop finetuning and CFG scheduling. Across all four datasets, ReDiFine successfully generates more “reusable” images, which can be reused to finetune a diffusion model without causing a severe model collapse (Section 5).
@@ 2 Related Work
The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [15, 10, 5, 16, 14, 12, 3, 7, 13, 4, 8, 9]. Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [6], has been observed in both language and image generative models.
Empirical studies on LLMs [13] reveal that linguistic diversity collapses, especially in high-entropy tasks [12], although this can be mitigated with data accumulation [2]. In image generation, several works [15, 10, 16, 17, 14, 3] note image degradation when diffusion models are recursively trained with self-generated data. We conduct extensive empirical experiments to identify the most significant factor causing model collapse in text-to-image diffusion models. Our findings reveal that while model collapse is universally observed across various datasets and scenarios, it manifests in three distinct types of image degradation.
@@
Figure 1: (a) Overall pipeline of the Chain of Diffusion. Given a pretrained text-to-image diffusion model $M_{0}$ and a prompt set $P$, a finetuned model $M_{k}$ is trained using $D_{k-1}$ generated at the previous iteration $k-1$. Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$, building a fully synthetic loop. Chain of diffusion begins with the original real dataset $D_{0}$. (b) Image degradation universally occurs during Chain of Diffusion across four datasets. As the Chain of Diffusion progresses, the severity of image degradation intensifies, exhibiting universal patterns of highly saturated images containing repetitive high-frequency patterns. This consistently holds across four datasets and $10$ scenarios that we comprehensively investigate in Section 3.2.
Theoretical studies on model collapse have largely focused on the reduction of diversity, typically framed as either decreasing covariance in continuous domains [16, 6, 3] or shrinking support in discrete domains [11, 9]. The finite number of generated samples per iteration has been identified as a key cause of model collapse by several authors [3, 6, 22, 11, 8], while others [16, 23, 9] emphasize that sampling bias reduces the generative model’s effective distribution. In contrast, we present a novel theoretical framework based on quantitative trait modeling, shifting the focus from variance reduction to mean drift induced by the selection process. We further demonstrate that many existing theories of model collapse can be understood as extensions of classical results from statistical genetics.
Regarding mitigation strategies for model collapse, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [16, 3, 8, 23] or accumulating additional data [2]. The only mitigation strategy beyond altering the training data composition is proposed by [4], who suggests a self-correcting self-consuming loop using an expert model to correct synthetic outputs. While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not be feasible and could be too costly for many real-world applications. In our work, we propose an alternative solution through reusable image generation, and show that it can mitigate model collapse more effectively than mixing more real data.
3 Model Collapse in Self-Consuming Chain of Diffusion
@@ 3.1 Problem Setting & Experimental Setup
Chain of Diffusion.
We begin with formally defining the self-consuming Chain of Diffusion finetuning. Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$. $M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a fully synthetic loop [16] for finetuning a pretrained generative model. Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
@@
During the Chain of Diffusion, $M_{0}$ and $P$ are fixed across all the iterations. To maintain the same size of training dataset, we generate one image per text prompt for all iterations. The overall pipeline of the Chain of Diffusion is shown in Figure 11(a).
Figure 2: Left: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse. All experiments were conducted on Pokemon except for the dataset size. For dataset size, we use CelebA since its original dataset is bigger, and we can subsample $500$, $1000$, and $2000$ images. For prompt, we concatenate prompts with different lengths.44footnotemark: 4 Right: All hyperparameter settings other than changing CFG show a high collapse rate greater than $100$, i.e., FID score increases by $\sim$2x in 6 iterations, indicating severe image degradation. The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the collapse rate defined in Equation (3). For both $\text{FID}_{1}$ and collapse rate, lower is better.55footnotemark: 5
Model and datasets.
We use Stable Diffusion v1.5 [24] as the pretrained model $M_{0}$ and apply LoRA [25] to finetune the Stable Diffusion at each iteration. We build our implementation on [26] and perform experiments on four datasets: Pokemon [27], Kumapi [28], Butterfly [29], and CelebA-1k [30] to investigate various domains including animation, handwriting, and real pictures. During each iteration, we finetune the pretrained Stable Diffusion $M_{0}$ for 100 epochs. More implementation details can be found in Appendix A.1 and A.2.
Evaluation metrics.
We use Frechet Inception Distance (FID) [31] to measure image fidelity. Following [32], we use DiNOv2 [33] as a feature extractor for FID since it is more consistent with our visual inspection than Inception-V3 network [34]. With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 features as the FID score. Quantitative results for additional metrics including CLIP score [35] can be found in Appendix E.
In addition to evaluation metrics for generative models presented above, we propose a new metric to quantify the reusability of images generated in the Chain of Diffusion. We define collapse rate as the performance degeneration per iteration in the Chain of Diffusion:
@@
where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set. We use FID as a performance metric here, but this can be CLIP, or any other performance metric of interest. Note that a low collapse rate indicates more reusable images since it means that the model does not degrade much during the Chain of Diffusion and we have $K=6$ in the rest of the paper.
@@ 3.2 Model Collapse in the Chain of Diffusion
Observation 1: Model collapse is universal in the Chain of Diffusion.
@@
We observe significant image degradation in all four datasets (see Figure 11(b)) in the Chain of Diffusion. The quality of generated images begins to clearly deteriorate in the third iteration and drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations. We investigated a variety of different scenarios (summarized in Figure 2) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon. We tested various dataset sizes for $D_{0}$ and $D_{k}$, increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt, mixing real images from $D_{0}$ to $D_{k}$, changing the descriptiveness of prompts, freezing U-Net or text encoder during finetuning, and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip). We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability, and investigated if the degradation happens for larger Stable Diffusion. In all settings we tested, image degradation was universally present and very fast. We plot the results for Pokemon on Figure 2 where y-axis is collapse rate and x-axis is the ${\rm FID}_{1}$ (better when closer to the origin). While adding noise to images and mixing $90\%$ real images to every iteration as proposed by [16, 3, 8, 23, 2] show the lowest collapse rate, they still exhibit significant degradation as shown in quantitative values (FID score has been doubled in $6$ iterations). Visual inspection for images is provided in Appendix C.
Observation 2: CFG is the most significant factor that impacts the model collapse.
@@
Throughout all our experiments, classifier-free guidance (CFG) had the biggest impact in the speed of model collapse. CFG scale was first introduced in [36] to modulate the strength between unconditional and conditional scores at each diffusion step as follows:
@@
A higher CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images. On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images. For those familiar with temperature sampling [37], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
@@
In Figure 2, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$ on x-axis), which is expected from our understanding of CFG. A more surprising part is that this comes at the cost of a worse collapse rate (an increase on the y-axis). Additionally, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and collapse rate. Similarly, when the CFG scale is too low—below $2.0$—the improvement in the collapse rate plateaus and both ${\rm FID}_{1}$ and collapse rate begin to increase. There is an optimal region of CFG values (near $2.5$, specific to Pokemon), where we achieve low collapse rate while maintaining a good quality in the first iteration as well. Moreover, Figure 33(c) presents FID for Pokemon to demonstrate how different CFG scales affect the performance during iterations. CFG scale $2.5$ achieves the most robust performance for Pokemon across iterations. Interestingly, optimal CFG scales differ for different styles: $2.5$ for animated or hand-writing datasets (Pokemon and Kumapi) and $1.5$ for photo datasets (CelebA and Butterfly) as shown in Appendix B.
(a) Degradations across different CFGs.
(b) t-SNE of images.
(c) FID $\downarrow$.
Figure 3: (a) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chain of Diffusion. CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down. More images in Appendix B. (b) t-SNE plot visualizes how generated images evolve from the original distribution (black) and shows distinct paths for three CFG scales (Pokemon). Different CFG scales and iterations are differentiated with different colors and transparency. Arrows indicate how the distributions of generated images move for different CFG scales. While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (7.5 and 1.0) deviate fast, indicating image degradation. (c) Quantitative comparison for FID $\downarrow$ (Pokemon). CFG $2.5$ achieves the most robust performance. CFG $1.0$ degrades from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in (a).
Observation 3: High CFG scales cause high-frequency degradation and low CFG scales cause low-frequency degradation.
@@
The CFG scale affects affect not only the speed of model collapse, but also the pattern of model collapse. As shown in Figure 33(a), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to as low-frequency degradation. On the other hand, for CFG $7.5$ the degradation pattern looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated. The t-SNE plot in Figure 33(b) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales. These patterns were consistent in all four datasets and detailed results can be found in Appendix B. In Section 4, we detail how different patterns of power spectra in high-frequency regions from different CFG scales can be understood using the framework from genetic biology.
@@
Implications of our observations. Our extensive investigations reveal that a high CFG of $7.5$, a common choice to generate visually appealing images, significantly increases collapse rate to achieve slightly better $\rm{FID}_{1}$. Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [16]. While they reported a monotonic increase in collapse rate as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$. It shows an intriguing trade-off between perceptual quality and reusability. This suggests that developers concerning reusability of images can substantially improve future generations by carefully choosing CFG.

### v3

@@ Abstract
Model collapse, the severe degradation of generative models when iteratively trained on their own outputs, has gained significant attention in recent years. This paper examines Chain of Diffusion, where a pretrained text-to-image diffusion model is finetuned on its own generated images. We demonstrate that severe image quality degradation was universal and identify CFG scale as the key factor impacting this model collapse. Drawing on an analogy between the Chain of Diffusion and biological evolution, we then introduce a novel theoretical analysis based on quantitative trait modeling from statistical genetics. Our theoretical analysis aligns with empirical observations of the generated images in the Chain of Diffusion. Finally, we propose Reusable Diffusion Finetuning (ReDiFine), a simple yet effective strategy inspired by genetic mutations. It operates robustly across various scenarios without requiring any hyperparameter tuning, making it a plug-and-play solution for reusable image generation.
@@ 1 Introduction
Can state-of-the-art AI models learn from their own outputs and improve themselves? As generative AI models (e.g., GPT, Diffusion) now churn out uncountable synthetic texts and images, this question piqued curiosity from many researchers in the past couple of years. While some show positive results of self-improving [1, 2], most report an undesirable “model collapse”—a phenomenon where a model’s performance degrades when it goes through multiple cycles of training with the self-generated data [3, 4, 5, 6, 7, 8, 9, 10]. When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text that has a lot of repetitions [11], and its linguistic diversity declines rapidly [12, 13]; image models also show quality degradation [14, 15] and loss of diversity [16, 17].
@@
The goal of this paper is to investigate model collapse in the practical scenario of finetuning pretrained text-to-image diffusion models. End users often finetune the latest model to generate images with a specific style (e.g., creating characters in the style of Pokémon). In fact, hundreds of new finetuned models trained on different styles and tons of generated images are regularly uploaded on platforms like CivitAI [^1] https://civitai.com/. When users scrape the Internet to collect images in the style they want, it becomes almost inevitable that synthetic images will be included in their datasets.
@@
To this end, we conduct a thorough empirical investigation into how various hyperparameters commonly used during diffusion finetuning (e.g., learning rate, diffusion steps, prompts) impact model collapse. We make a crucial observation: the classifier-free guidance (CFG) scale is the most impactful factor affecting the rate of model collapse. Moreover, we observe a fascinating phenomenon that the direction of degradation varies—low CFG results in low-frequency degradation in images, while high CFG leads to high-frequency degradation. The critical role of CFG in both the rate and direction of model collapse is a novel insight previously unrecognized in the literature.
@@
To gain a deeper theoretical understanding, we draw upon the concept of quantitative trait modeling from statistical genetics. Unlike existing studies that attribute model collapse to limited sample sizes and conclude it leads to zero variance [3, 6, 16], we show that a truncation-based selection results in mean drift of quantitative features. We found the high-frequency power spectra of generated images as one of the features that accurately follow our theoretical modeling. Moreover, we show that other theoretical work on model collapse can also be connected to statistical genetics models, such as random genetic drift and its variants [18, 19, 20, 21]. This fresh angle of drawing parallels with statistical genetics offers a new framework to understand the mechanism of model collapse.
While we have shown the crucial role of CFG in mitigating model collapse both empirically and theoretically, searching for the optimal CFG scale over iterations is computationally expensive and cumbersome to implement in practice. We propose a simple solution that bypasses CFG tuning, called Reusable Diffusion Finetuning (ReDiFine). It introduces small modifications in the finetuning and generation—condition drop and CFG scheduling—and we demonstrate that ReDiFine achieves performance comparable to the optimal CFG scale across all hyperparameter settings we tested.
Our contributions can be summarized as follow:
  • We conducted a comprehensive investigation of model collapse when finetuning a diffusion model on its own outputs, testing a wide range of parameters on four datasets (two digital art and two natural images). Our results show that CFG scale is the most critical factor, governing both the rate of model collapse and the type of image degradation (Section 3).
@@
  • We propose ReDiFine, a simple yet effective strategy to achieve a near-optimal reusability-fidelity trade-off without any hyperparameter tuning. By combining condition drop finetuning and CFG scheduling, ReDiFine operates robustly across various scenarios—such as synthetic-real mixed datasets and varying initial CFG settings—across all four datasets we evaluated (Section 5).
@@ 2 Related Work
The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [3, 4, 5, 7, 8, 9, 10, 12, 13, 14, 15, 16]. Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [6], has been observed in both language and image generative models [^2] The definition of model collapse is expanding [22], but we focus on a fully self-consuming training loop [16]..
Empirical studies on LLMs [13] reveal that linguistic diversity collapses, especially in high-entropy tasks [12], although this can be mitigated with data accumulation [2]. In image generation, several works [3, 10, 14, 15, 16, 17] note image degradation when diffusion models are recursively trained with self-generated data. We conduct extensive experiments to reveal the key factor causing model collapse in text-to-image diffusion models. Our findings reveal that while model collapse is universally observed across various datasets and scenarios, it manifests in three distinct types of image degradation.
Theoretical studies on model collapse largely focused on diversity reduction, typically framed as either decreasing covariance in continuous domains [3, 16, 23] or shrinking support in discrete domains [9, 11], which aligns with our observations on decreased recall. The finite number of generated samples has been identified as a key cause of model collapse [3, 8, 11, 23], while others [9, 16, 24] emphasize that sampling bias reduces the effective distribution. There also exists the study about recursive stability of generative models in the self-consuming training loops [25]. In contrast, we present a novel framework based on quantitative trait modeling, providing a fresh direction based on mean drift induced by the selection process. We further demonstrate that many existing theories of model collapse can be understood through extensions of classical statistical genetics.
Most existing works echo the importance of incorporating a large proportion of real data [3, 8, 16, 24, 25] or accumulating data over iterations [2] to mitigate model collapse. [4] instead suggests a self-correcting self-consuming loop using an expert model, a physics simulator for human motion generation, to correct synthetic outputs. Several concurrent works seek to address model collapse through their theoretical insights. [26] studies the effects of synthetic data in scaling laws, showing how the scales of models affect model collapse behaviors. Moreover, [27] and [28] emphasize the importance of data verification in preventing performance degradation, while [29] introduces a novel token editing strategy for data curation in LLM tasks. On the other hand, [30] and [31] propose model-specific solutions leveraging real data to counteract the negative effects of synthetic data in image-generative models. In our work, we propose an alternative solution through reusable image generation, and show that it can generate high-quality images comparable to the optimal hyperparameter without extensive search.
3 Model Collapse in Self-Consuming Chain of Diffusion Finetuning
@@
Figure 1: (a) Overall pipeline of the Chain of Diffusion. Given a pretrained text-to-image diffusion model $M_{0}$ and a prompt set $P$, a finetuned model $M_{k}$ is trained using $D_{k-1}$ generated at the previous iteration $k-1$. Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$, building a fully synthetic loop. Chain of diffusion begins with the original real dataset $D_{0}$. (b) Image degradation universally occurs across multiple domains, and low and high CFG scales lead to blurry images and high-frequency degradation, respectively. As the Chain of Diffusion progresses, the severity of image degradation intensifies, which holds consistently across four datasets and $11$ scenarios in Section 3.2. CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down. More images in Appendix C.
@@ 3.1 Problem setting & Experimental Setup
Chain of Diffusion. We begin with formally defining the self-consuming Chain of Diffusion finetuning. Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$. $M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a fully synthetic loop [16]. Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
@@
During the Chain of Diffusion, $M_{0}$ and $P$ are fixed across iterations, and one image is generated per one prompt to maintain the dataset size for all iterations. The overall pipeline is shown in Figure 1(a). We note that this setting is consistent with [6, 8, 12, 13, 16].
Model and datasets. We use Stable Diffusion v1.5 [32] as the pretrained model $M_{0}$ and apply LoRA [33] to finetune $M_{0}$ at each iteration. We build our implementation on [34] and experiment on four datasets: Pokemon [35], Kumapi [36], Butterfly [37], and CelebA-1k [38] to investigate various domains including animation, handwriting, and real pictures. $M_{0}$ is finetuned for 100 epochs during each iteration. More details can be found in Appendix B.1 and B.2.
Evaluation metrics. We use Frechet Inception Distance (FID) [39] to measure image quality. Following [22], we use DiNOv2 [40] as a feature extractor since it is more consistent with our visual inspection than Inception-V3 network [41]. With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 as the FID score. Additional experiments with other metrics (CLIP score [42], Sample-wise Feature Distance (SFD), and recall) are presented in Appendix F.5.
We propose a new metric to quantify the degradation of generated images. We define collapse rate as the performance degeneration per iteration in the Chain of Diffusion:
@@
where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set. We use FID as a performance metric here, but this can be any other performance metric of interest. Note that a low collapse rate indicates more reusable images since the model does not degrade much. We have $K=6$ for the rest of the paper.
Figure 2: Left: Description of $11$ potential factors (excluding CFG) that we examined as candidate sources for model collapse. All experiments were conducted on Pokemon except for the dataset size. For dataset size, we use CelebA since its original dataset is bigger, and we can subsample $500$, $1000$, and $2000$ images. For prompt, we concatenate prompts with different lengths. Center: All hyperparameter settings other than changing CFG show a high collapse rate greater than 1.0, i.e., FID score increases by $\sim$2x in 6 iterations, indicating severe image degradation. The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the collapse rate defined in Eqn. 3. For both $\text{FID}_{1}$ and collapse rate, lower is better. Right: Quantitative comparison for FID $\downarrow$ (Pokemon). CFG $2.5$ achieves the most robust performance. CFG $1.0$ degrades from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in Figure 1(b).
@@ 3.2 Model Collapse in the Chain of Diffusion
Observation 1: Model collapse is universal in the Chain of Diffusion. We observe significant image degradation in all four datasets in the Chain of Diffusion (see Figure 1(b) and Appendix C). The quality begins to deteriorate in the third iteration and drops rapidly once the degradation emerges, reaching an unrecognizable level in two or three additional iterations. Quantitative evaluation (FID, CLIP score, SFD, and recall) also indicates this degradation (Appendix F.5). We observed significant drops in recall in all experiments, indicating diversity reduction discussed in several recent works.
@@
We then investigated a variety of different scenarios (summarized in Figure 2) to see if this degradation is an anomaly of specific hyperparameters or a ubiquitous phenomenon. We tested various dataset sizes for $D_{0}$ and $D_{k}$, increasing the size of $D_{k}$ by generating more than one image per prompt, mixing real images of $D_{0}$ to $D_{k}$, changing the descriptiveness of prompts [^3] C, W, S, and L for Combine, Waifu, Short, and Long, respectively. We concatenate BLIP and Waifu captions as default setting, referred to as Combine. More details can be found in Appendix D.4., freezing U-Net or text encoder, and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip). We also added a small Gaussian noise in each image in the original set $D_{0}$ to test if small random perturbations can improve reusability, or merged all four datasets into one to see if the extended domain delays model collapse (Cross-domain data). In all settings we tested, image degradation was universally present and very fast. We plot the trade-off for Pokemon on Figure 2 where y-axis is collapse rate and x-axis is the ${\rm FID}_{1}$ (better when closer to the origin). While adding noise to images and mixing $90\%$ real images to every iteration as proposed by [2, 3, 8, 16, 24] show the lowest collapse rate, they still exhibit significant degradation (FID score has been doubled in $6$ iterations) [^4] We do not display scenarios that change the training dataset size, such as Img/Prompt and Dataset Size, as varying sizes differ FID scales. We observe similar image degradation for them, presented in the Appendix D..
Moreover, this degradation is unavoidable for Stable Diffusion XL (Appendix D.11). We further examined different settings from the literature where all previously generated images are accumulated (data accumulation [2], Appendix D.12) or the same model is iteratively finetuned (iterative retraining [3, 15], Appendix F.4), and consistently observed model collapse.
(a) t-SNE of images.
(b) Selection mechanism.
(c) Power spectra of images vs. Mean of phenotypes.
Figure 3: (a) t-SNE plot visualizes how generated images evolve from the original distribution (black) and shows distinct paths for three CFG scales (Pokemon). Different CFG scales and iterations are differentiated with different colors and transparency. Arrows indicate how the distributions of generated images move for different CFG scales. While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (1.0 and 7.5) deviate fast, indicating image degradation. (b) Selection mechanism with two-sided truncation. $r_{1}$ and $r_{2}$ ratios of samples are truncated from the left and right tails, and the remaining $r=1-r_{1}-r_{2}$ ratio of samples is selected. (c) Power spectra for different CFGs (Pokemon) and quantitative trait modeling results align well for different selection strategies. Directional selections with truncation can effectively explain our observations in Section 3: the behaviors of high- and low-frequency degradation and optimal CFG scale. Detailed settings are provided in Appendix E.3.
@@
Observation 2: CFG is the most significant factor that impacts the model collapse. Throughout all our experiments, classifier-free guidance (CFG) had the biggest impact on the speed of model collapse. CFG scale was first introduced in [43] to modulate the balance between unconditional score and conditional score at each diffusion step as follows:
Total Score=Unconditional Score
+CFG(Conditional ScoreUnconditional Score)
Total ScoreUnconditional Score
CFGConditional ScoreUnconditional Score
@@
High CFG emphasizes the conditional score for the given prompt, which pushes the generation to align better with the prompt and often leads to higher-fidelity images. On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images. For those familiar with temperature sampling [44], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
@@
Figure 2 shows that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$ on x-axis), which is expected from our understanding of CFG. Surprisingly, this comes at the cost of a worse collapse rate (an increase on the y-axis). When the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and it worsens both ${\rm FID}_{1}$ and collapse rate. Similarly, when the CFG scale is too low—below $2.0$—the improvement in collapse rate plateaus, and both ${\rm FID}_{1}$ and collapse rate begin to increase. There is an optimal region of CFG values (near $2.5$, specific to Pokemon), where we achieve a low collapse rate while maintaining low ${\rm FID}_{1}$ as well. Figure 2 presents FID for Pokemon to demonstrate how different CFG scales affect performance. CFG scale $2.5$ achieves the most robust performance for all metrics. Interestingly, optimal CFG scales differ for different styles: $2.5$ for animated or hand-writing datasets (Pokemon and Kumapi) and $1.5$ for photo datasets (CelebA and Butterfly) as shown in Appendix C.
@@
Observation 3: High CFG scales cause high-frequency degradation and low CFG scales cause low-frequency degradation. CFG scale does not only affect the speed of model collapse, but also the pattern of model collapse. As shown in Figure 1(b), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to low-frequency degradation. On the other hand, for CFG $7.5$ how images degrade looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated. The t-SNE plot in Figure 3(a) clearly demonstrates distinct paths for high, low, and medium CFG scales over iterations. They were consistent in all four datasets (Appendix C). In Section 4, we detail how different patterns of high-frequency power spectra from different CFG scales can be understood using the framework from genetic biology.
@@
Implications of our observations. We showed that a high CFG of $7.5$, a common choice to generate visually appealing images, significantly increases collapse rate to achieve slightly better $\rm{FID}_{1}$. Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [16]. While they reported a monotonic increase in collapse rate as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not monotonic when we look CFG scales from $1.0$ to $10.0$. It shows an intriguing trade-off between perceptual quality and reusability, suggesting developers to consider reusability of images to achieve substantially improved future generations by carefully choosing CFG.